Vitex, Inc. manufactures a popular consumer product and it has provided the following data excerpts from its standard cost system:

Inputs (1) Standard Quantity or Hours (2)
Standard
Price
or Rate
Standard
Cost
(1) × (2)
Direct materials 2.20 pounds $ 16.20 per pound $ 35.64
Direct labor 1.00 hours $ 15.40 per hour $ 15.40
Variable manufacturing overhead 1.00 hours $ 9.50 per hour $ 9.50
Total standard cost per unit $ 60.54
Total Variances Reported
Standard
Cost*
Price
or Rate
Quantity or
Efficiency
Direct materials $ 570,240 $ 10,788 F $ 32,400 U
Direct labor $ 246,400 $ 3,400 U $ 15,400 U
Variable manufacturing overhead $ 152,000 $ 4,800 F $ ? U

*Applied to Work in Process during the period.

The company's manufacturing overhead cost is applied to production on the basis of direct labor-hours. All of the materials purchased during the period were used in production. Work in process inventories are insignificant and can be ignored.

Required:

1. How many units were produced last period?

2. How many pounds of direct material were purchased and used in production?

3. What was the actual cost per pound of material? (Round your answer to 2 decimal places.)

4. How many actual direct labor-hours were worked during the period?

5. What was the actual rate paid per direct labor-hour? (Round your answer to 2 decimal places.)

6. How much actual variable manufacturing overhead cost was incurred during the period?
Answer #1

1) No of unit produced = 570240/35.64 = 16000 Units

2) Direct material quantity variance = (Standard qty-actual qty)Standard price

-32400 = (16000*2.2*16.20-16.20X)

16.20X = 570240+32400

X(actual quantity) = 37200

3) Material price variance = (Standard price-actual price)actual quantity

10788 = (16.20*37200-37200X)

37200X = 591852

X(actual price) = 15.91

4) Labor efficiency variance = (Standard hour-actual hour)Standard rate

-15400 = (16000*15.40-15.40X)

15.40X = 261800

X(Actual hour) = 17000

5) Labor rate variance = (Standard rate-actual rate)actual hour

-3400 = (15.40*17000-17000X)

17000X = 265200

X(actual rate) = 15.60

6) Variable overhead efficiency variance = (16000-17000)*9.5 = 9500 U

Actual overhead = 152000-4800+9500 = 156700